mary has 18 coins, some quarters and the rest dimes. the total value of all the coins is $3.45. Find the number of each of coin.

Answer by stanbon(57282)

mary has 18 coins, some quarters and the rest dimes. the total value of all the coins is $3.45. Find the number of each of coin.
----------------------------------------------------
Equations:
Quantity Equation: q + d = 18
Value Equation...:25q+10d=345 cents
----------------------------------------
Multiply thru the 1st equation by 25 to solve for "d":
25q + 25d = 25*18
25q + 10d = 345
-------------------------
Subtract the 2nd equation from the 1s to get:
15d = 25*18 - 345
d = 7 (# of dimes)
---
Substitute into q + d = 18 to solve for "q":
q + 7 = 18
q = 11 (# of quarters)
====================================
Cheers,
Stan H.